OpenStudy (anonymous):

Given the equation 6x − 1 = 11, which order of operations completely solves for x?

OpenStudy (anonymous):

6x-1=11 6x=11+1 6x=12 x=2

OpenStudy (anonymous):

Add 1, then divide by 6 Add 11, then divide by −1 Subtract 1, then divide by 6 Subtract 11, then divide by −1

OpenStudy (anonymous):

these are the answers

OpenStudy (anonymous):

add 1 then divide by 6

OpenStudy (skullpatrol):

apply the inverse operations in the opposite order

OpenStudy (anonymous):

A function is created to represent the miles per gallon your car gets. What restrictions would be made to the domain? @skullpatrol

OpenStudy (anonymous):

The domain would only include integers. The domain would only include positive integers. The domain would include all real numbers. The domain would only include positive numbers.

OpenStudy (anonymous):

the domain would only include positive numbers

OpenStudy (anonymous):

thanks